OpenStudy (anonymous):

Help Please!! You live in a state with a sales tax of 5%. You buy a car for $15,000. What is the sales tax? a) $50 b) $500 c) $750 d) $1,500

jimthompson5910 (jim_thompson5910):

5% = 5/100 = 0.05 multiply 0.05 with the price of the car to get the sales tax

jimthompson5910 (jim_thompson5910):

Rule: To convert a percent to a decimal, divide by 100 \[\Large x \% = \frac{x}{100}\]

OpenStudy (anonymous):

thank you so much! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem, glad to be of help